What is the boiling point of a solution that contains 12.5g biphenyl, C12H10, in 100g of bromobenzene? The normal boiling point of bromobenzene is 156°C, and kb for bromobenzene is +6.26°C/m.


1
Expert's answer
2021-10-22T03:57:34-0400

∆ T = mKb

Molar mass of bromobenzene = 157.02 g/mol

100/157.02= 0.64mol

Molar mass of biphenyl= 154.21

12.5/154.21= 0.081mol

0.64+0.081= 0.721×6.26= 4.51

= 4.51+156= 160.51°C